OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80

OpenStudy (amistre64):

if its geometric, then youll need to determine the common ratio

OpenStudy (amistre64):

in general: a1 = a1 a2 = a1*r a3 = a2*r , or a1*r*r, or a1*r^2 a4 = a3*r, or a1*r^3 a5 = a1*r^4 ... an = a1*r^(n-1)

OpenStudy (amistre64):

notice that since a2 = a1*r we can algebra this to: r = a2/a1

OpenStudy (anonymous):

-5x(-4)^n-1